# RE D shift: quasinormal modes of black holes embedded in halos of matter

### v1

@@ IV.1 A result of sufficient generality concerning light ring & QNM frequencies in the eikonal regime
Before discussing QNMs, we will show that for any sufficiently dilute halo of matter, the light ring frequency and Lyapunov timescale are redshifted, by an amount that scales with the halo compactness. Since light ring properties are intimately connected to QNMs [96], this result then provides a robust basis to expect that to leading order QNM frequencies will be redshifted by environments.
@@
Notice that for a mass $M$ distributed on a scale $a_{0}\gg M_{\rm BH}$ then $U_{\rm LR}\sim U(0)\sim-M/a_{0}$, and indeed this coincides with the large $a_{0}$ limit of the exact solution (4). The potential $U_{\rm LR}$ gives the redshift of a photon as it crosses the galaxy, due to its matter content. The above – a key result of this work which explains our numerics below – then indicates that the leading order consequence of matter distributions is a redshift relative to vacuum modes, as had also been observed via time-domain signals [64].
@@ IV.2 Axial QNMs of BHs within Hernquist and NFW matter profiles
Our main numerical results are summarized in Figs. 25. Figure 2 shows the real and the imaginary part of the fundamental and first overtone for Hernquist-type BHs, with varying $\ell$ and halo compactness $M/a_{0}$, where $(\alpha,\beta,\gamma)=(1,4,1)$. As expected, increasing $\ell$ yields QNM frequencies with larger real component, approaching the angular frequency of null particles at the light ring [96]. Our results show that more compact halo configurations lead to QNMs with lower (in absolute value) real and imaginary part, regardless of the overtone number and of the mode index $\ell$.
@@
The accuracy of the relation (21) improves for realistic environments with $M/a_{0}\leq 10^{-4}$. This confirms results obtained for the exact BH metric (2), for which QNMs have been computed using a hyperboloidal foliation scheme [97, 98, 99, 100, 42, 101] and a spectral collocation method [102, 103, 104, 105, 106, 107, 108], proving the accuracy of our generalized matrix method [96, 25]. The lower panel of Fig. 3 also shows the relative difference between frequencies computed with the WKB approach, which differ, at most, less than $1\%$ from the generalized matrix method values.
@@
where we have introduced the redshift factor $\gamma\equiv 1-M/a_{0}$ and [83]:
@@ V Conclusions
In fact, since almost all BHs reside in some sort of environment (accretion disks, galactic and extra-galactic) the effort towards a complete BH spectroscopy program that includes environmental effects is undergoing rapid advancements. Very recently, a fully-relativistic, static and spherically symmetric BH was found that is embedded in a spherical matter halo [67]. This solution does not contain impromptu post-Newtonian additions to the metric that mimic environmental effects. To the contrary, it is a fully-relativistic BH solution of the Einstein field equations surrounded by an anisotropic generic fluid.
@@
Understanding the environmental effect on BH QNMs is now very timely and of utmost importance for current and future detections of GWs. In fact, a spurt of analyses have taken place very recently for a unique, exact, BH solution surrounded by a Hernquist-type cold matter [67, 68, 87, 55]. Since this solution is the only one that currently exist in closed form (besides some thin-shell matching solutions [69] and analytic approximations [70]), using different matter galactic halos leads to numerical BH solutions surrounded by these matter halos [72, 75].
@@
Here, we have initiated the use of numerically-constructed BH configurations with different matter halos to find their axial QNMs and observe the overall effect of the environment to the modes. We show that BHs with environments possess scalar, electromagnetic and axial gravitational QNMs that are governed by a single-parameter effective potential. The analyses in [67, 25] have shown that the exact solution from Ref. [67] possesses scalar and axial QNMs that are redshifted, both in frequency and (absolute value of) decay timescale. In this work, we solidify that this effect on axial QNMs is universal, no matter the energy density (in our study, numerical Hernquist and numerical NFW-type BHs). This is shown by the use of matrix method from Ref. [88, 89] which we generalized so that it works for numerically-constructed backgrounds. The code is publicly available in Ref. [109] and can be used not only for environmental effects, but also for other kinds of solutions such as hairy BHs that do not possess an exact, closed form, and BHs with accretion disks.
@@
Overall, our code behaves very well, in accord to all tests we have performed, and produces convergent QNMs (see Appendix B). Since axial perturbations do not couple to the surrounding matter, a redshift of modes, as compactness increases, is adequate. Therefore, the difference between axial QNMs due to environmental effects can be equivalent to redshifted vacuum Schwarzschild QNMs. However, we expect redshift scaling to break for polar perturbations, which couple to matter variables [75]. The polar sector of gravitational perturbations couples to matter and this will enable us to test GR in its strongest regime. Calculations of polar QNMs for different families of matter profiles are underway and will be addressed in a followup publication, along with an assessment of their detectability by future GW observations.
@@
Note: During the completion of our work, Ref. [110] appeared, that contains results for axial QNMs, for an exact metric, that are partially comparable to ours.
@@ Appendix A Generalized matrix method
One of the numerical computation techniques to solve eigenvalue problems of differential equations is the matrix method [88, 89]. With this method, it is possible to transform the Schrödinger-like equation into a matrix equation via the proper manipulation of the ordinary differential equation that governs perturbation propagation through proper decomposition of the spatial derivatives.
The proposed matrix method in [88, 89] is a non-grid-based interpolation approach, where the interpolation is based on Taylor expansions. This method discretizes the spatial coordinate so that the differential equation, as well as its boundary conditions, are transformed into a homogeneous matrix equation. With this procedure the eigenfunction is expanded in the vicinity of all grid points, and therefore the precision of the algorithm is potentially improved. The data points do not need to sit on a specified grid for the method to work (equispaced in [89], delimited expansion [92] and Chebyshev points [93]). Using the information of a set of $N$ randomly scattered data points, Taylor series are carried out for the unknown eigenfunction up to $N$-th order for each data point. Then, the resulting homogeneous system of linear algebraic equations is solved for the eigenvalues.
@@
By using the Cramer’s rule, it is possible to discretize all derivatives [88, 89] and rewrite (18) in the matrix form
@@ Appendix B QNM convergence
Finally, Fig. 7 demonstrates the convergence of the fundamental axial QNM for the Herquist and NFW-type BHs for different $M/a_{0}$. We assume that the modes resulting from the generalized matrix method with $N=19$ are accurate enough to be the reference modes and show that as $N$ increases then the axial $n=0$ QNMs indeed converge to those found and utilized throughout the paper. We observe that the percentage error gradually decreases as $N$ is increased from 3 to 17. The convergence is rather fast, especially when compared to other methods that use discretization techniques such as pseudospectral [102] and hyperboloidal methods [97], where a grid point number of $N=50$ is needed to achieve the same accuracy as the one shown in Fig. 7.

### v2

@@ IV.1 A result of sufficient generality concerning light ring & QNM frequencies in the eikonal regime
Before discussing QNMs, we will show that for any sufficiently dilute halo of matter, the light ring frequency and Lyapunov timescale are redshifted, by an amount that scales with the halo compactness. Since light ring properties are intimately connected to QNMs Cardoso et al. (2009), this result then provides a robust basis to expect that to leading order QNM frequencies will be redshifted by environments.
@@
Notice that for a mass $M$ distributed on a scale $a_{0}\gg M_{\rm BH}$ then $U_{\rm LR}\sim U(0)\sim-M/a_{0}$, and indeed this coincides with the large $a_{0}$ limit of the exact solution (4). The potential $U_{\rm LR}$ gives the redshift of a photon as it crosses the galaxy, due to its matter content. The above – a key result of this work which explains our numerics below – then indicates that the leading order consequence of matter distributions is a redshift relative to vacuum modes, as had also been observed via time-domain signals Spieksma et al. (2024).
@@ IV.2 Axial QNMs of BHs within Hernquist and NFW matter profiles
Our main numerical results are summarized in Figs. 36. Figure 3 shows the real and the imaginary part of the fundamental and first overtone for Hernquist-type BHs, with varying $\ell$ and halo compactness $M/a_{0}$, where $(\alpha,\beta,\gamma)=(1,4,1)$. As expected, increasing $\ell$ yields QNM frequencies with larger real component, approaching the angular frequency of null particles at the light ring Cardoso et al. (2009). Our results show that more compact halo configurations lead to QNMs with lower (in absolute value) real and imaginary part, regardless of the overtone number and of the mode index $\ell$.
@@
The accuracy of the relation (21) improves for realistic environments with $M/a_{0}\leq 10^{-4}$. This confirms results obtained for the exact BH metric (2), for which QNMs have been computed using a hyperboloidal foliation scheme Ansorg and Panosso Macedo (2016); Panosso Macedo et al. (2018); Panosso Macedo (2020); Hennig and Panosso Macedo (2021); Jaramillo et al. (2021); Panosso Macedo and Zenginoglu (2024) and a spectral collocation method Jansen (2017); Cardoso et al. (2018a, b); Destounis (2019a); Liu et al. (2019); Destounis (2019b); Destounis et al. (2019), proving the accuracy of our generalized matrix method Cardoso et al. (2009); Berti et al. (2025). The lower panel of Fig. 4 also shows the relative difference between frequencies computed with the WKB approach, which differ, at most, less than $1\%$ from the generalized matrix method values.
@@
where we have introduced the redshift factor $\gamma\equiv 1-M/a_{0}$ and Regge and Wheeler (1957):
@@ V Conclusions
In fact, since almost all BHs reside in some sort of environment (accretion disks, galactic and extra-galactic) the effort towards a complete BH spectroscopy program that includes environmental effects is undergoing rapid advancements. Very recently, a fully-relativistic, static and spherically symmetric BH was found that is embedded in a spherical matter halo Cardoso et al. (2022a). This solution does not contain impromptu post-Newtonian additions to the metric that mimic environmental effects. To the contrary, it is a fully-relativistic BH solution of the Einstein field equations surrounded by an anisotropic generic fluid.
@@
Understanding the environmental effect on BH QNMs is now very timely and of utmost importance for current and future detections of GWs. In fact, a spurt of analyses have taken place very recently for a unique, exact, BH solution surrounded by a Hernquist-type cold matter Cardoso et al. (2022a); Konoplya (2021); Cardoso et al. (2022b, 2024). Since this solution is the only one that currently exist in closed form (besides some thin-shell matching solutions Jusufi (2023) and analytic approximations Konoplya and Zhidenko (2022)), using different matter galactic halos leads to numerical BH solutions surrounded by these matter halos Figueiredo et al. (2023); Speeney et al. (2024).
@@
Here, we have initiated the use of numerically-constructed BH configurations with different matter halos to find their axial QNMs and observe the overall effect of the environment to the modes. We show that BHs with environments possess scalar, electromagnetic and axial gravitational QNMs that are governed by a single-parameter effective potential. The analyses in Cardoso et al. (2022a); Berti et al. (2025) have shown that the exact solution from Ref. Cardoso et al. (2022a) possesses scalar and axial QNMs that are redshifted, both in frequency and (absolute value of) decay timescale. In this work, we solidify that this effect on axial QNMs is universal, no matter the energy density (in our study, numerical Hernquist and numerical NFW-type BHs). This is shown by the use of matrix method from Ref. Lin and Qian (2016, 2017) which we generalized so that it works for numerically-constructed backgrounds. The code is publicly available in Ref. git and can be used not only for environmental effects, but also for other kinds of solutions such as hairy BHs that do not possess an exact, closed form, and BHs with accretion disks.
@@
Overall, our code behaves very well, in accord to all tests we have performed, and produces convergent QNMs (see Appendix B). Since axial perturbations do not couple to the surrounding matter, a redshift of modes, as compactness increases, is adequate. Therefore, the difference between axial QNMs due to environmental effects can be equivalent to redshifted vacuum Schwarzschild QNMs. However, we expect redshift scaling to break for polar perturbations, which couple to matter variables Speeney et al. (2024). The polar sector of gravitational perturbations couples to matter and this will enable us to test GR in its strongest regime. Calculations of polar QNMs for different families of matter profiles are underway and will be addressed in a followup publication, along with an assessment of their detectability by future GW observations.
@@
Note: During the completion of our work, Ref. Chakraborty et al. (2024) appeared, that contains results for axial QNMs, for an exact metric, that are partially comparable to ours.
@@ Appendix A Generalized matrix method
One of the numerical computation techniques to solve eigenvalue problems of differential equations is the matrix method Lin and Qian (2016, 2017). With this method, it is possible to transform the Schrödinger-like equation into a matrix equation via the proper manipulation of the ordinary differential equation that governs perturbation propagation through proper decomposition of the spatial derivatives.
The proposed matrix method in Lin and Qian (2016, 2017) is a non-grid-based interpolation approach, where the interpolation is based on Taylor expansions. This method discretizes the spatial coordinate so that the differential equation, as well as its boundary conditions, are transformed into a homogeneous matrix equation. With this procedure the eigenfunction is expanded in the vicinity of all grid points, and therefore the precision of the algorithm is potentially improved. The data points do not need to sit on a specified grid for the method to work (equispaced in Lin and Qian (2017), delimited expansion Lin and Qian (2023) and Chebyshev points Shen et al. (2023)). Using the information of a set of $N$ randomly scattered data points, Taylor series are carried out for the unknown eigenfunction up to $N$-th order for each data point. Then, the resulting homogeneous system of linear algebraic equations is solved for the eigenvalues.
@@
By using the Cramer’s rule, it is possible to discretize all derivatives Lin and Qian (2016, 2017) and rewrite (18) in the matrix form
@@ Appendix B QNM convergence
Finally, Fig. 8 demonstrates the convergence of the fundamental axial QNM for the Herquist and NFW-type BHs for different $M/a_{0}$. We assume that the modes resulting from the generalized matrix method with $N=19$ are accurate enough to be the reference modes and show that as $N$ increases then the axial $n=0$ QNMs indeed converge to those found and utilized throughout the paper. We observe that the percentage error gradually decreases as $N$ is increased from 3 to 17. The convergence is rather fast, especially when compared to other methods that use discretization techniques such as pseudospectral Jansen (2017) and hyperboloidal methods Ansorg and Panosso Macedo (2016), where a grid point number of $N=50$ is needed to achieve the same accuracy as the one shown in Fig. 8.